TheEmotinalOne:

A surveyor asked 250 residents of a town with a population of 15,000 on whether they fish in any of the town's lakes. 95 people answered "yes" to the question. How many people would be expected to answer "yes" given the entire population was surveyed? Enter your answer in the box. people

Sherixn:

So try to find the percent of people who said yes to the question, and then using that percent, see how many people in the entire population would say yes.

Sherixn:

If you need any more help, feel free to ask <3

jhonyy9:

like an idea use the rule of 3 simple and in this way you can write from 250 asked answered yes 95 people from 15000 asked answered x ,cross multiplie and calcule the x x *250 = 95 *15000 x = 95 * 15000 / 250 = ? hope helped understandably easy